# The Power of Graph Sparsification in the Continual Release Model

### v1

@@ 3 Preliminaries
We introduce the notation we use in this paper as well as some standard privacy and graph theoretic definitions. The familiar reader can feel free to skip this section.
@@ 3.1 Setting & Notation
We consider the insertion-only setting within the continual release model, where we begin with an empty graph $G_{0}=(V,E_{0})$ where $E_{0}=\varnothing$ and edge updates in the form of $\{e_{t},\perp\}$ arrive in a stream at every timestep $t\in[T]$. We are required to release an output for the problem of interest after every edge update. Our goal is to obtain algorithms that achieve sublinear space, either in the number of edges $\tilde{o}(m)$ (note that $T\geq m$ as we allow for empty updates), or the number of vertices $\tilde{o}(n)$. We assume $T\leq n^{c}$ for some absolute constant $c\geq 1$ to simplify our presentation. If there are no empty edge updates then it is sufficient to consider $c=2$.
@@
3.2 Approximation Algorithms
We consider optimization problems in both the minimization and maximization setting. Let $\operatorname{OPT}\in\mathbb{R}$ be the optimum value for the problem. For a minimization problem, a $(\beta,\zeta)$-approximate algorithm outputs a solution with cost at most $\beta\cdot\operatorname{OPT}+\zeta$. For a maximization problem, $(\beta,\zeta)$-approximate algorithm outputs a solution of value at least $\frac{1}{\beta}\cdot\operatorname{OPT}-\zeta$. For estimation problems, a $(\beta,\zeta)$-approximate algorithm outputs an estimate $\tilde{R}\in\mathbb{R}$ of some quantity $R\in\mathbb{R}$ where $R-\zeta\leq\tilde{R}\leq\beta\cdot R+\zeta$ (one-sided multiplicative error). As a shorthand, we write $\beta$-approximation algorithm to indicate a $(\beta,0)$-approximation algorithm.
3.3 Concentration Inequalities
Below, we make use of the multiplicative Chernoff bound.
Theorem 3.1 (Multiplicative Chernoff Bound; Theorems 4.4, 4.5 in [mitzenmacher2015probability]).
  1. 1.
    Upper Tail: $\mathbf{P}(X\geq(1+\eta)\cdot\mu)\leq\exp\left(-\frac{\eta^{2}\mu}{2+\eta}\right)$ for all $\eta>0$;
  2. 2.
    Lower Tail: $\mathbf{P}(X\leq(1-\eta)\cdot\mu)\leq\exp\left(-\frac{\eta^{2}\mu}{3}\right)$ for all $10<\eta<$ 1.
3.4 Neighboring Streams
@@
We use $G_{t}$ and $E_{t}$ to denote the graph induced by the set of set of updates in the stream $S$ up to and including update $t$. Now, we define neighboring streams as follows. Intuitively, two graph streams are node-neighbors if one can be obtained from the other by removing all the edge updates incident to a single vertex. (For graph streams, the edges adjacent to the vertex are spread over many timesteps). Similarly, two graph streams are edge neighbors if one can be obtained from the other by removing one edge update (replacing the edge update by an empty update in a single timestep).
@@
Definition 3.4 (Node Neighboring Streams).
Two streams of edge updates, $S=[u_{1},\dots,u_{T}]$ and $S^{\prime}=[u^{\prime}_{1},\dots,u^{\prime}_{T}]$, are node-neighboring if there exists exactly one vertex $v^{*}\in V$ where for all $t\in[T]$, $u_{t}\neq u^{\prime}_{t}$ only if $u_{t}$ or $u_{t}^{\prime}$ is an edge addition or deletion of an edge adjacent to $v^{*}$. Streams may contain any number of empty updates, i.e. $u_{t}=\bot$. Without loss of generality, we assume for the updates $u_{t^{*}}\neq u^{\prime}_{t^{*}}$ that $u^{\prime}_{t^{*}}=\bot$ and $u_{t^{*}}=\pm e_{t}$ is an edge insertion or deletion of an edge adjacent to $v^{*}$.
We make a short note of comparison between this definition to the definition of node-neighboring graphs in the static setting. While node-neighboring graphs in the static setting are defined in terms of the difference of all edges adjacent to any one node, node-neighboring streams are defined in terms of all updates adjacent to any one node. This means that for fully dynamic streams where both edge insertions and deletions can occur, a particular edge can be inserted and deleted many times; this entire sequence of insertions and deletions on the same edge could differ between node-neighboring streams. For insertion-only streams, such a distinction is moot since any edge can be associated with at most one update. Note that our edge-neighboring streams are defined in terms of any one update event so such a distinction between insertion-only and fully dynamic streams also does not exist for edge-neighboring streams.
We now formalize the concepts of edge edit distance between streams and the concept of distance-preserving.
Definition 3.5 (Edge/Node Edit Distance).
The node edit distance between two streams is the length of the shortest chain where every adjacent pair of streams in the chain are node-neighboring.
An algorithm that takes as input a stream $S$ and outputs a chosen set of updates from the stream is distance-preserving if on edge-neighboring streams $S$ and $S^{\prime}$, there exists a coupling between the randomness used in the algorithm on the inputs such that the edge distance between the output streams is $O(1)$. We say such an algorithm is distance-preserving on node-neighboring streams if given two node-neighboring streams $S$ and $S^{\prime}$, there exists a coupling such that the output streams have edge edit distance $O(\tilde{\alpha})$ where $\tilde{\alpha}$ is some public parameter passed into the algorithm.
3.5 Differential Privacy Tools
Here, we define the privacy tools commonly used in differential privacy in terms of the continual release model.
@@ 3.5 Differential Privacy Tools
Definition 3.7 (Node Differential Privacy).
Let $\varepsilon,\delta\in(0,1)$. An algorithm $\mathcal{A}(S):\mathcal{S}^{T}\rightarrow\mathcal{Y}^{T}$ that takes as input a graph stream $S\in\mathcal{S}^{T}$ is said to be $(\varepsilon,\delta)$-node differentially private (DP) if for any pair of node-neighboring graph streams $S,S^{\prime}$, that differ in all updates adjacent to one vertex, and for every $T$-sized vector of outcomes $Y\subseteq\operatorname{Range}(\mathcal{A})$,
When $\delta=0$, we say that $\mathcal{A}$ is $\varepsilon$-node DP.
Throughout the paper, we use some standard privacy mechanisms as building blocks (see [dwork2014algorithmic] for a reference).
@@ 13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
4 Maximum Matching and Arboricity Sparsification
@@
We begin this section in Section 4.1 with an edge-DP algorithm for maximum matching that uses truly sublinear space. We also briefly sketch how to remove the assumption on $\tilde{\alpha}$ in Section E.2 at the cost of more space and a worse approximation guarantee.
Then, we derive space-efficient algorithms that satisfy the stronger node-differential privacy guarantee (cf. Definition 3.7) for graphs of bounded arboricity $\alpha$. The key idea is that a judicious choice of a sparsification algorithm reduces both the space complexity and the edge edit distance between node-neighboring graphs to some factor $\Lambda=O(\alpha)$. Such sparsification is useful since, in the worst case, the edge edit distance pre-sparsification can be $\Omega(n)$. Then, it suffices to run any edge-DP algorithm with privacy parameter $\nicefrac{{\varepsilon}}{{\Lambda}}$ after the sparsification to achieve the same privacy guarantee but reduced error.
By adapting techniques from [solomon2021local], we derive such sparsifiers for the maximum matching problem in Section 4.2 (and the minimum vertex cover problem in Section 5). Then, we design a node-DP maximum matching algorithm in the continual release setting in Section 4.3. Later, in Section 5, we explain how to release a node-DP implicit vertex cover in the semi-streaming setting.
4.1 Edge-DP Maximum Cardinality Matching
@@
In this section, we adapt the algorithm of [mcgregor2018simple] to obtain a private approximate maximum cardinality algorithm in the continual release model using small, sublinear space in the number of vertices. As in their algorithm, we assume that we are provided with a public upper bound $\tilde{\alpha}$ on the maximum arboricity $\alpha$ of the graph at any point in the stream. The privacy of our algorithm is always guaranteed. However, we do not assume that $\tilde{\alpha}$ is guaranteed to upper bound $\alpha$; when $\tilde{\alpha}\geq\alpha$, then our approximation guarantees hold with high probability. Otherwise, our approximation guarantees do not necessarily hold. Note that the same type of guarantee holds for the original non-private streaming algorithm [mcgregor2018simple] where their utility guarantee is only given when their public estimate of $\alpha$ upper bounds the maximum arboricity of the input. We prove the following result in this section.
@@
Fix $\eta\in(0,1]$. Given a public estimate $\tilde{\alpha}$ of the maximum arboricity $\alpha$ over the stream [^11] We can eliminate this assumption with an additional pass of the stream or notify the observer when the utility guarantees no longer hold in the one-pass setting. See Appendix E., Algorithm 4.1 is an $\varepsilon$-edge DP algorithm for estimating the size of the maximum matching in the continual release model for insertion-only streams. If $\tilde{\alpha}\geq\alpha$, then with probability at least $1-1/\operatorname*{poly}(n)$, our algorithm returns a $\left((1+\eta)(2+\tilde{\alpha}),O\left(\frac{\log^{2}(n)}{\varepsilon\eta}\right)\right)$-approximation of the size of the maximum matching at every timestamp. Moreover, our algorithm uses $O\left(\frac{\log^{2}(n)\log(\tilde{\alpha})}{\varepsilon\eta^{2}}\right)$ space with probability $1-1/\operatorname*{poly}(n)$.
4.1.1 Algorithm Description
Algorithm Intuition

### v3

@@ 4 Preliminaries
We now introduce the notation we use in this paper as well as some definitions. Further standard preliminaries are deferred to Appendix A.
@@ 4.1 Setting & Notation
We consider the insertion-only setting within the continual release model, where we begin with an empty graph $G_{0}=(V,E_{0})$ where $E_{0}=\varnothing$ and edge updates in the form of $\{e_{t},\perp\}$ arrive in a stream at every timestep $t\in[T]$ where each non-empty edge is inserted at most once. [^7] This is without loss of generality as redundant edge insertions can be treated as an empty update. We are required to release an output for the problem of interest after every (possibly empty) update. Our goal is to obtain algorithms that achieve sublinear space, either in the number of edges $\tilde{o}(m)$ (note that $T\geq m$ as we allow for empty updates), or the number of vertices $\tilde{o}(n)$. We assume $T\leq n^{c}$ for some absolute constant $c\geq 1$ to simplify our presentation. If there are no empty edge updates then it is sufficient to consider $c=2$.
@@
4.2 Continual Release
@@
We use $G_{t}$ and $E_{t}$ to denote the graph induced by the set of set of updates in the stream $S$ up to and including update $t$. Now, we define neighboring streams as follows. Intuitively, two graph streams are edge neighbors if one can be obtained from the other by removing one edge update (replacing the edge update by an empty update in a single timestep).
@@
The notion of neighboring streams leads to the following definition of an edge differentially private algorithm.
@@
We now formalize the concepts of edge edit distance between streams and the concept of stability under sparsification.
Definition 4.4 (Edge Edit Distance).
An algorithm that takes as input a stream $S$ and outputs a chosen set of updates from the stream is stable if on edge-neighboring streams $S$ and $S^{\prime}$, there exists a coupling888 A coupling of random variables $(X,Y)$ is a joint distribution such that the marginal distributions correspond to $X,Y$ respectively. between the randomness used in the algorithm on the inputs such that the edge distance between the output streams is $O(1)$.
4.3 Differential Privacy Tools
Here, we define the privacy tools commonly used in differential privacy in terms of the continual release model. Throughout the paper, we use some standard privacy mechanisms as building blocks (see [dwork2014algorithmic] for a reference).
@@ 13                   $\operatorname{count}\leftarrow\operatorname{count}+1$
5 $k$-Core Decomposition
In this section, we introduce a semi-streaming sampling algorithm that preserves the $k$-cores in an input graph $G=(V,E)$ while ensuring privacy. Specifically, we have the following theorem.
Theorem 5.1 (Sublinear Space Private $k$-Core Decomposition).
Fix $\eta\in(0,1]$. Algorithm B.2 is an $\varepsilon$-DP algorithm for the $k$-core decomposition problem in the continual release model for insertion-only streams. At every $t\in[T]$, the algorithm returns a value for each vertex, such that every value is a $\left(2+\eta,O\left({\eta^{-2}}\varepsilon^{-1}\log^{3}(n)\right)\right)$-approximation of the corresponding vertex’s true core value, with probability $1-1/\operatorname*{poly}(n)$. The maximum space used is $O\left({\eta^{-4}}\varepsilon^{-1}n\log^{5}n\right)$, with probability $1-1/\operatorname*{poly}(n)$.
Algorithm Intuition. First, we give some intuition for our algorithm (Algorithm B.2). Our algorithm essentially performs a sampling version of the classic peeling algorithm for $k$-core decomposition. The classic peeling algorithm successively peels (removes) vertices with the minimum degree until all vertices are removed from the graph. A core that is formed during the peeling process is the induced subgraph consisting of the remaining vertices after a vertex is peeled and the value of such a core is the minimum induced degree within the subgraph. The core number for each vertex $v$ is equal to the maximum valued core that $v$ is a part of during any stage of the peeling. A dynamic version of this algorithm can be obtained by maintaining a level data structure where a vertex is moved up a level if its induced degree among vertices in the same or higher levels is larger than a cutoff $C$. One can show that having $O(\log n)$ levels of the structure and appropriately setting $C$ among $O(\log n)$ duplicates of the structure gives a $(2+\eta)$-approximation of the core numbers of the nodes in the non-private, insertion-only setting [SCS20, LSYDS22, DLRSSY22]. Our main innovation is a private and sparse level data structure.
Finally, to make the above algorithm differentially private, we use SVT to determine when to move the vertex up a level. We show that although many vertices may move up levels, we only lose privacy when the vertices that are adjacent to the edge that differs between neighboring streams move up. Since our total number of levels and duplicates is bounded by $O(\log^{2}n)$, the privacy loss from SVT is also bounded by $O(\log^{2}n)$.
Analysis. We provide the pseudocode and proof of Theorem 5.1 in Appendix B. Specifically, the pseudocode of the sparse level data structure is provided in Algorithm B.1 and the pseudocode of the actual algorithm is in Algorithm B.2. Section B.1 then presents a detailed description of the algorithm and we prove its privacy and utility guarantees in Sections B.2 and B.3, respectively.
6 Densest Subgraph
Theorem 6.1 (Sublinear Space Private Densest Subgraph).
Fix $\eta\in(0,1]$. Algorithm C.2 is an $\varepsilon$-edge differentially private algorithm for the densest subgraph problem in the continual release model for insertion-only streams. The algorithm returns a set of vertices whose induced subgraph is a $\left(2+\eta,O({\eta^{-1}}\varepsilon^{-1}\log^{2}(n))\right)$-approximation of the densest subgraph in $G_{t}$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$. The maximum space used is $O\left({\eta^{-2}}\varepsilon^{-1}n\log^{2}(n)\right)$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$.
We can also reduce the multiplicative error to $(1+\eta)$ at the cost of increasing the space usage by a $\operatorname*{poly}(\log(n))$ factor.
Theorem 6.2 (Sublinear Space Private Densest Subgraph).
Fix $\eta\in(0,1]$. There exists an $\varepsilon$-edge differentially private algorithm for the densest subgraph problem in the continual release model for insertion-only streams. The algorithm returns a set of vertices whose induced subgraph is a $\left(1+\eta,{\eta^{-4}}\varepsilon^{-1}\log^{5}(n)\right)$-approximation of the densest subgraph in $G_{t}$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$. The maximum space used is $O\left({\eta^{-5}}\varepsilon^{-1}n\log^{5}(n)\right)$, with probability at least $1-1/\operatorname*{poly}(n)$, for all $t\in[T]$.
Algorithm Intuition. We revise the algorithms of [mcgregor2015densest] and [EHW16] to the insertion-only continual release setting. On a high-level, our algorithm maintains a sample of edges over time and releases a DP set of vertices by running a black-box DP densest subgraph algorithm (e.g., Theorem F.2) on the subgraph induced by the sample. At every timestep $t\in[T]$, an edge $e_{t}$ is sampled with probability $p$ — the sampling probability is initialized to 1 as in the beginning we can afford to store every edge. We privately check whether the number of edges seen so far exceeds a certain threshold using a sparse vector technique (SVT) query and adjust the sampling probability $p$ accordingly. In order to avoid privacy loss that grows linearly in $T$, we do not invoke the black-box DP densest subgraph algorithm at every timestep and instead invoke it only if the current density of the sample exceeds a certain threshold using another SVT query.
Analysis. The pseudocode and proof of Theorems 6.1 and 6.2 are deferred to Appendix C. In particular, pseudocode for a useful data structure is provided in Algorithm C.1 and the main algorithm is in Algorithm C.2. Section C.1 describes the algorithm in detail and we prove its privacy and utility guarantees in Sections C.2 and C.3, respectively.
7 Maximum Matching
@@
In this section we give an edge-DP algorithm for maximum matching that uses truly sublinear space. We adapt the algorithm of [mcgregor2018simple] to obtain a private approximate maximum cardinality algorithm in the continual release model using sublinear space in the number of vertices. As in their algorithm, we assume that we are provided with a public upper bound $\tilde{\alpha}$ on the maximum arboricity $\alpha$ of the graph at any point in the stream. We also briefly sketch how to remove the assumption on $\tilde{\alpha}$ in Section H.2 at the cost of more space and a worse approximation guarantee.
@@
The privacy of our algorithm is always guaranteed. However, we do not assume that $\tilde{\alpha}$ is guaranteed to upper bound $\alpha$. When $\tilde{\alpha}\geq\alpha$, our approximation guarantees hold with high probability. Otherwise, our approximation guarantees do not necessarily hold. Note that the same type of guarantee holds for the original non-private streaming algorithm [mcgregor2018simple] where their utility guarantee is only given when their public estimate of $\alpha$ upper bounds the maximum arboricity of the input. We prove the following result in this section.
@@
Fix $\eta\in(0,1]$. Given a public estimate $\tilde{\alpha}$ of the maximum arboricity $\alpha$ over the stream, [^9] We can eliminate this assumption with an additional pass of the stream or notify the observer when the utility guarantees no longer hold in the one-pass setting. See Appendix H. Algorithm D.1 is an $\varepsilon$-edge DP algorithm for estimating the size of the maximum matching in the continual release model for insertion-only streams. If $\tilde{\alpha}\geq\alpha$, then with probability at least $1-1/\operatorname*{poly}(n)$, our algorithm returns a $\left((1+\eta)(2+\tilde{\alpha}),O(\eta^{-1}\varepsilon^{-1}\log^{2}(n))\right)$-approximation of the size of the maximum matching at every timestamp. Moreover, our algorithm uses $O(\eta^{-2}\varepsilon^{-1}{\log^{2}(n)\log(\tilde{\alpha})})$ space with probability $1-1/\operatorname*{poly}(n)$.